OpenStudy (anonymous):

y = (x – 2)^2(x – 1) y = 3x^3 – 3x Find the zeroes?

hartnn (hartnn):

to find the zeros, put y= 0 in each of them

OpenStudy (anonymous):

So y = (x – 2)^2(x – 1) (x-2)^2(x-1)=0 I don't know what to do next for this one. y = 3x^3 – 3x 3x^3-3x=0 This one divide by 3 x^3-x=0 Then I'm stuck again

hartnn (hartnn):

ok, when we have \((x-a)(x-b) =0 \) we can say that either \(x-a = 0 \quad or \quad x-b =0 \)

OpenStudy (anonymous):

Yeah I got that.

hartnn (hartnn):

so if we have \((x-2)^2(x-1) = 0 \) what can we say ?

OpenStudy (anonymous):

(x-2)^2=0; x=? The square throws me off. (x-1)=0; x=1

hartnn (hartnn):

ok, (x-a)^2 = 0 means x-a = 0 or x-a = 0 so, x = a or a which means x= a with the multiplicity of 2 !

hartnn (hartnn):

is that new to you ?

OpenStudy (anonymous):

I'm still a little confused So I have (x-2)^2=0 x-2=0 or x-2=0? They'd be the same x=2 according to what you wrote that's how I should do it.

hartnn (hartnn):

yes, same roots or EQUAL roots instead of saying, x = 2,2 we say that x= 2 with the multiplicity of 2!

OpenStudy (anonymous):

Oh ok that kind of makes sense. So my answer would be x=1, 2 multiplicity of 2

hartnn (hartnn):

yes, thats correct :)

OpenStudy (anonymous):

Okay great!

hartnn (hartnn):

for x^3-x = 0 first factor out x from the left

OpenStudy (anonymous):

So x(x^2-1)=0?

hartnn (hartnn):

correct! so, x= 0 or x^2 -1 =0 can you solve x^2 -1 =0 ??

hartnn (hartnn):

its easy :)

OpenStudy (anonymous):

(x-1)(x+1)=0 x-1=0 x+1=0 x=0,1,-1

hartnn (hartnn):

correct!

hartnn (hartnn):

good work :)

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

most welcome ^_^